OpenStudy (anonymous):

What is the value of the expression –5x4th power when x = –2? A. 80 B. 40 C. –40 D. –80 What is the value of the expression –5 times-2 to the 4th power? i have done this problem over and over and my answer has been 80 repeatedly. but when i submitted my test it says its -80.

jimthompson5910 (jim_thompson5910):

Keep in mind that x^2 is always positive even if x is negative example: x = -5 x^2 = x*x = (-5)*(-5) = +25

jimthompson5910 (jim_thompson5910):

through similar logic, x^4 is always positive example: x = -3 x^4 = x*x*x*x = (-3)*(-3)*(-3)*(-3) = +81

jimthompson5910 (jim_thompson5910):

I'm using the idea (negative) * (negative) = positive

jimthompson5910 (jim_thompson5910):

so if x^4 is always positive, then -5x^4 is always negative because (negative) * (positive) = negative

jimthompson5910 (jim_thompson5910):

does that help @nanchita150 ?

OpenStudy (anonymous):

thank you @jim_thompson5910 it is letting me respond now and i figured the medal thing out as well

jimthompson5910 (jim_thompson5910):

you're welcome